Problem A.1


The graph below is made of three line segments:


-1 1 2 3 4 5 6 7 8 9 10 11 12


1


2


3


4


y


x


f(x)


g(x)


h(x)


The segments correspond to the following three functions:


f(x) = x − 2, g(x) = p


4 − (x − 6)2 + 2, h(x) = x − 6


Find the total length L of the graph between x = 2 and x = 10.



Problem A.2


Let f(x), g(x) and h(x) be the functions from Problem A.1. Find the derivative λ




(x) of the


following function with respect to x:


λ(x) = f(x) · g(x) + f(x) · h(x) − g(x) · h(x)



*Please give specific answers to both Problem A.1 & A.2


1
Expert's answer
2021-11-24T13:49:57-0500

A 1




"f(x) = x \u2212 2" for "x\isin [2,4]"

"g(x) = \sqrt{4 − (x − 6)^2} + 2" for "x\isin [4,8]"

"h(x) = x \u2212 6" for "x\isin [8,10]"


then:

length of f(x):

"L_1=\sqrt{2^2+2^2}=2\sqrt 2"


length of h(x):

"L_2=\sqrt{2^2+2^2}=2\sqrt 2"


length of g(x):

"L_3=2\pi R/2=2\pi"


total length:

"L=L_1+L_2+L_3=2\sqrt 2+2\sqrt 2+2\pi=2(\pi+2\sqrt 2)"


A 2


"f(x) · g(x)=(x-2)(\sqrt{4 − (x − 6)^2} + 2)"

"f(x) \u00b7 h(x)=(x-2)(x-6)"

"g(x) · h(x)=(x-6)(\sqrt{4 − (x − 6)^2} + 2)"

"λ(x)=(x-2)(\sqrt{4 − (x − 6)^2} + 2)+(x-2)(x-6)-(x-6)(\sqrt{4 − (x − 6)^2} + 2)"


"λ'(x)=\sqrt{4 − (x − 6)^2} + 2-\frac{(x-2)(x-6)}{\sqrt{4 − (x − 6)^2} }+2x-8+\sqrt{4 − (x − 6)^2} + 2-\frac{(x-6)^2}{\sqrt{4 − (x − 6)^2} }="


"=2\sqrt{4 − (x − 6)^2}+2x-4-\frac{(2x-8)(x-6)}{\sqrt{4 − (x − 6)^2} }"